OpenStudy (anonymous):

solve by factoring or completing the square: 4x^2+30=29x

OpenStudy (anonymous):

\[(x-6) (4 x-5) = 0\]

OpenStudy (anonymous):

which one is that? factoring or completing the square?

Directrix (directrix):

That is factoring after rewriting 4x^2+30=29x as 4x^2 - 29x + 30 = 0. The roots are 6 and 5/4.

OpenStudy (anonymous):

 completing the square